OpenStudy (anonymous):

Please somebody explain this to me 1/4 + 1/7 x 1/9

OpenStudy (anonymous):

\[\frac{1}{4}+\frac{1}{63}=\frac{67}{252} \]

OpenStudy (anonymous):

How did you get the answer though? please expalin thank you for your help

OpenStudy (anonymous):

( 1* 63+4*1)/(4*63)

OpenStudy (anonymous):

Thank you 100% you have really helped me understand it...thank you your explaination was so much better than my book..your great!!!